# A coupled-trajectory approach for decoherence, frustrated hops and internal consistency in surface hopping

### v2

@@ Abstract
We address the issues of decoherence, frustrated hops and internal consistency in surface hopping. We demonstrate that moving away from an independent-trajectory picture is the strategy which allows us to propose a robust surface hopping scheme overcoming all these issues at once. Based on the exact factorization and on the idea of coupled trajectories, we consider the swarm of trajectories, that mimic the nuclear dynamics in nonadiabatic processes, as a unique entity. In this way, imposing energy conservation of the swarm and allowing the trajectories to share energy when hops occur clearly indicates the route towards a new surface hopping scheme. Encouraging results are reported, in terms of electronic and vibrational time-dependent properties on the photodynamics of fulvene, modeled with a full-dimensional linear vibronic coupling Hamiltonian.
@@
Nonadiabatic molecular dynamics is an irreplaceable method in the toolkit of physical chemists to model time-resolved photochemistry, the study of the coupled electron-nuclear evolution of molecules upon light excitation[1]. Among the many theories available nowadays, surface hopping, since its first appearance in the literature in 1971[2], has grown to become the method of choice to simulate the nonadiabatic dynamics of molecules and to accompany experimental observations.[3, 4, 5, 6, 7, 8, 9, 10, 11] Despite alternative theories being continuously developed, surface hopping remains very popular thanks to the unrivaled accuracy-cost compromise, to its relative implementation simplicity and especially to the fact that it conveys a very intuitive picture of a nonadiabatic process.[12, 13, 14, 15, 16, 17] In surface hopping, the nuclear wavepacket is represented in terms of a swarm of independent trajectories that classically evolve on the support of electronic potential energy surfaces obtained as the solution of the time-independent Schrödinger equation at a given nuclear configuration. The key nonadiabatic effects, beyond the Born-Oppenheimer approximation, i.e., transitions between electronic states, are mimicked by stochastic hops of the trajectories from one potential energy surface to another usually occurring in localized regions of configuration space. Each trajectory of the swarm is coupled to the quantum-mechanical evolution of the electronic time-dependent state – even though not all surface hopping implementations require this.[18, 19]
@@
Surface hopping has its weaknesses, which have been reported extensively in the literature: it suffers from overcoherence, due to the disconnect between the dynamics of the classical nuclei and of the quantum electrons[20, 21, 22, 23, 24, 25, 16, 26, 27, 28]; when hops occur, energy conservation for the hopping trajectory is enforced, which causes frustrated hops if a trajectory hopping to a state of higher potential energy does not have enough kinetic energy to compensate for the increase in potential[29]; velocity rescaling related to the energy conservation after the hops is applied quite arbitrarily and strongly influence the dynamics;[30, 31, 32, 33, 34] using a swarm of independent trajectories to mimic the evolution of a quantum wavepacket misses non-local quantum-mechanical effects[22, 35, 19, 36, 37]. These weaknesses are consequences of the fact that surface hopping was not deduced originally from first principles, even though derivations [38, 35, 39] and qualitative justifications [40, 41] were proposed later on.
In this work, we address the question: What is missing in surface hopping that would cure all its weaknesses at once? We demonstrate that the idea of coupled trajectories emanating from the exact factorization [42, 43, 44, 45] is capable to indicate the route to follow. Specifically, this route implies that the “surface hopping” idea is used as a tool to simplify conceptually and numerically the evolutions equations derived from the exact factorization, rather than as the working framework. In this way, we do not need to cure the weaknesses of surface hopping, since the starting point is a formulation which is naturally free from these weaknesses. Therefore, we propose a theoretical framework that fully exploits the coupling of the trajectories and that is combined with the algorithmic efficiency of surface hopping, bypassing introducing the issues of decoherence, frustrated hops and internal consistency. Note that a coupled-trajectory-based surface hopping scheme has been developed previously [19, 36], but, as we will show below, it inherits some of the issues of the original surface hopping since the idea of coupled trajectories was not fully “embraced” by all aspects of the algorithm. Instead, our new proposition is robust, stable and accurate, and, thus, we believe that our strategy is an important step forward in the development of novel ideas for nonadiabatic molecular dynamics schemes.
@@
The idea of coupled trajectories emerges naturally in the framework of the exact factorization (EF).[42, 43] When the time-dependent molecular wavefunction is factored as the product of a marginal nuclear amplitude, or wavefunction, and a conditional electronic amplitude, the molecular time-dependent Schrödinger equation (TDSE) can be decomposed as an effective nuclear TDSE and an electronic evolution equation. The nuclear dynamics is generated by time-dependent vector and scalar potentials, reminiscent of classical electromagnetic potentials, representing the effect of the electrons. Formally, an elegant and rigorous procedure can be introduced to derived classical-like Hamilton’s equations from a TDSE, which is the procedure followed in context of EF to deduce classical-like nuclear trajectories.[46] In turn, the nuclei affect the electronic dynamics entering the evolution equation with their momentum/velocity field and with their spatial distribution. As simple as this, the coupled-trajectory idea emerges in the exact factorization: while the momentum field can be interpreted classically, as it is a local property that can be encoded in a trajectory, the very nature of the spatial nuclear distribution is non local. Therefore, to determine the quantum momentum, as we dubbed the term depending on the nuclear distribution, the trajectories become coupled since only the whole ensemble yields the necessary non-local information. Extensive analyses of the time-dependent potentials of the EF have justified additional approximations, namely (i) the use of purely adiabatic forces to drive the nuclear dynamics, since the time-dependent scalar potential reproduces piecewise the adiabatic shapes [40], and (ii) the idea of hopping with associated velocity rescaling, to compensate for the effect of the time-dependent vector potential that couples to the nuclear kinetic energy in the nuclear TDSE [41, 47]. Based on these observations, the coupled-trajectory Tully surface hopping (CT-TSH) algorithm was proposed [19, 36].
@@
CT-TSH is capable to circumvent the issue of overcoherence via the presence of the quantum momentum in the electronic evolution, which encodes, as argued above, information about the nuclear spatial delocalization. Nonetheless, as we will show below and as previously observed [48], CT-TSH suffers from frustrated hops since, apart from the evaluation of the quantum momentum, the trajectories behave as independent entities and their coupling is not fully exploited, in particular to enforce energy conservation after the hops. Instead, by simply understanding that the trajectories need to cooperate to yield a meaningful approximation of the nuclear quantum dynamics, the additional issues discussed above can be easily cured: if the trajectories share their kinetic energy when hops occur, so as to conserve the total energy of the swarm, frustrated hops are avoided (or strongly reduced to a statistically irrelevant amount) and different velocity rescaling schemes do not affect the outcome of the simulation.
Understanding the importance of fully exploiting the coupling among the trajectories in surface hopping leads us to the derivation of the combined coupled-trajectory Tully surface hopping (CCT-TSH), that we illustrate and test in the remainder of this paper. CCT-TSH is based on the idea of “energy sharing” to conserve the energy of the swarm of trajectories and to alleviate the frustrated hops. We propose three schemes for the energy sharing, which are schematically summarized in Fig. 1: (i) the equity-based scheme, (ii) the overlap-based scheme, and (ii) the quantum-momentum-based scheme.
@@
Before detailing the energy-sharing schemes, let us introduce the following quantities. The trajectories are labeled with Greek indices, and the hopping trajectory is the trajectory $\alpha$, with kinetic energy indicated with the symbol $E_{\text{kin}}^{\alpha}$. The variation in potential energy for a hop from state $I$ towards a state $J$ of higher energy is $\Delta E^{\alpha}_{\text{hop}}<0$. The Cartesian components of the vectors are indicated with Latin indices, thus for instance the symbol $v_{i}^{\beta}(t)$ stands for the component $i$ of the velocity of the trajectory $\beta$ at time $t$. For scheme (ii), we define the overlap between two trajectories, i.e., $S_{\alpha\beta}$, as the overlap between two Gaussians centered at the positions of the trajectories $\alpha$ and $\beta$. Note that the algorithm does not depend on the value of the width of the Gaussians, as long as they are the same for all trajectories, since $S_{\alpha\beta}$ is only used to quantify the proximity of the trajectories. For scheme (iii), we identify two channels for energy sharing, one driven by the nonadiabatic couplings between states $I$ and $J$, $\boldsymbol{d}_{IJ}^{\alpha}$, and one driven by the quantum momentum, $\boldsymbol{Q}^{\alpha}$, both evaluated at the position of the hopping trajectory. Therefore, we use the scalar quantities $\boldsymbol{d}_{IJ}^{\alpha}\cdot\boldsymbol{v}^{\alpha}$ and $\boldsymbol{Q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}$, to decompose the energy variation to be compensated for the hop to occur. Note that these scalar quantities appear in the electronic evolution equation of CT-TSH and CCT-TSH, with $\boldsymbol{f}_{IJ}^{\alpha}$ the time-integrated difference of the adiabatic forces for states $I$ and $J$.
Using the definitions given above, the energy-sharing schemes can be summarized as follow:
@@
    where ${v_{i}^{\beta}}^{\prime}$ is the velocity after rescaling. Weighting the contribution of the trajectories $\beta$ depending on their kinetic energy ensures a minimal impact on the trajectory itself. Since the trajectory $\alpha$ uses its entire kinetic energy for the hop, its new velocity after the hop is $\boldsymbol{v}^{\alpha}=0$.
@@
    Quantum-momentum-based. The amount of energy to be rescaled $\Delta E^{\alpha}_{\text{hop}}$ is partitioned in two contributions: One contribution, $\Delta E^{\alpha}_{\text{NAC}}$, that will be compensated by the kinetic energy of the hopping trajectory $\alpha$ and that will be rescaled along the nonadiabatic coupling vector, and a second contribution, $\Delta E^{\alpha}_{\text{Qmom}}$, that will be compensated by the kinetic energy of the whole swarm of trajectories based on (ii). The two contributions are calculated according to
    $\Delta E^{\alpha}_{\text{NAC}}=\Delta E^{\alpha}_{\text{hop}}\frac{\boldsymbol{d}_{IJ}^{\alpha}\cdot\boldsymbol{v}^{\alpha}}{\boldsymbol{d}_{IJ}^{\alpha}\cdot\boldsymbol{v}^{\alpha}+\boldsymbol{Q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}}$ (3)

### v3

# A coupled-trajectory strategy for decoherence, frustrated hops and internal consistency in surface hopping

@@ Abstract
We address the issues of decoherence, frustrated hops and internal consistency in surface hopping. We demonstrate that moving away from an independent-trajectory picture is the strategy which allows us to propose a robust surface hopping scheme overcoming all these issues at once. Based on the exact factorization and on the idea of coupled trajectories, we consider the swarm of trajectories, that mimics the nuclear dynamics in nonadiabatic processes, as a unique entity. In this way, imposing energy conservation of the swarm and allowing the trajectories to share energy when hops occur clearly indicates the route towards a new surface hopping scheme. Encouraging results are reported, in terms of electronic and vibrational time-dependent properties on the photodynamics of fulvene and 4-(dimethyloamino)benzonitrile, modeled with a full-dimensional linear vibronic coupling Hamiltonian.
keywords:
American Chemical Society,
[…]
{tocentry}[Uncaptioned image]
1 Introduction
@@
Nonadiabatic molecular dynamics is an irreplaceable method in the toolkit of physical chemists to model time-resolved photochemistry, the study of the coupled electron-nuclear evolution of molecules upon light excitation1. Among the many theories available nowadays, surface hopping, since its first appearance in the literature in 19712, has grown to become the method of choice to simulate the nonadiabatic dynamics of molecules and to accompany experimental observations.3, 4, 5, 6, 7, 8, 9, 10, 11 Despite alternative theories being continuously developed, surface hopping remains very popular thanks to the unrivaled accuracy-cost compromise, to its relative implementation simplicity, and especially to the fact that it conveys a very intuitive picture of a nonadiabatic process.12, 13, 14, 15, 16, 17 In surface hopping, the nuclear densities associated with the electronic states participating in the photochemical reaction are represented in terms of swarms of independent trajectories that classically evolve on the support of the corresponding electronic potential energy surfaces. Such adiabatic, or Born-Oppenheimer, surfaces are calculated as the eigenvalues of the electronic time-independent Schrödinger equation at given nuclear configurations. The key nonadiabatic effects, beyond the Born-Oppenheimer approximation, i.e., transitions between electronic states, are mimicked by stochastic hops of the trajectories from one potential energy surface to another, usually occurring in localized regions of configuration space. Each trajectory of the swarm is coupled to the quantum-mechanical evolution of the electronic time-dependent state – even though not all surface hopping implementations require this.18, 19
@@
Surface hopping has weaknesses, which have been reported extensively in the literature: it suffers from overcoherence, due to the disconnect between the dynamics of the classical nuclei and of the quantum electrons20, 21, 22, 23, 24, 25, 16, 26, 27, 28; when hops occur, energy conservation for the hopping trajectory is enforced, which causes frustrated hops if a trajectory hopping to a state of higher potential energy does not have enough kinetic energy to compensate for the increase in potential29; the velocity rescaling method applied to enforce energy conservation after the hops can significantly affect the system’s dynamics;30, 31, 32, 33, 34, 35 using a swarm of independent trajectories to mimic the evolution of a quantum wavepacket misses non-local quantum-mechanical effects22, 36, 19, 37, 38. These weaknesses are consequences of the fact that surface hopping was not deduced from first principles, even though derivations 39, 36, 40 and qualitative justifications 41, 42 were proposed later on. Despite these weaknesses, as said above, surface hopping is a very powerful tool to model time-resolved photochemistry, as a recent Prediction Challenge 43 demonstrated. Outside of its “comfort zone”, however, surface hopping can perform poorly in its standard formulation, and care must be taken to assess possible extensions either of the theory or of the applications, for instance when dealing with external time-dependent fields, 44, 45 like laser pulses, and with the creation of electronic wavepackets. 46. In the literature, these situations are more and more often studied with nonadiabatic dynamics simulations as the fields of femtochemistry and attochemistry continue to evolve on the experimental side, 47, 48 which motivates the interest in continuing exploring alternative approaches to simulate nonadiabatic molecular dynamics.
Curing or alleviating the weaknesses of surface hopping remains a challenging and interesting problem from the theoretical viewpoint, especially if grounded on a rigorous quantum-mechanical basis, which might open the possibility of achieving a robust algorithm for extensions and generalizations. It is with this perspective that surface-hopping-based techniques continue to appear in the literature, as for instance consensus surface hopping (CSH), 36 surface hopping based on the exact factorization (SH-XF), 49 coupled-trajectory Tully surface hopping (CT-TSH), 19 the mapping approach to surface hopping (MASH), 50 quantum-trajectory surface hopping based on the exact factorization (QTSH-XF), 51 quantum-trajectory surface hopping with simplified decay of mixing (QTSH-SDM). 52 Motivated by these ongoing efforts and the clear interest of the physical-chemistry community in testing and benchmarking 53 nonadiabatic molecular dynamics methods, we propose here some ideas based on the exact factorization 54, 55, 56, 57 to tackle the weaknesses of surface hopping. Specifically, we aim to address the following question: What is missing in surface hopping that would cure all its weaknesses at once? We believe that the idea of coupled trajectories emanating from the exact factorization is capable of indicating the route to follow. Specifically, this route implies that the “surface hopping” scheme is only introduced as a tool to simplify conceptually and numerically the evolution equations derived from the exact factorization, while the fundamental issues responsible for decoherence – and the resulting internal inconsistency – are bypassed by the coupling among the trajectories. Note that a coupled-trajectory-based surface hopping scheme has been developed previously 19, 37, but, as we will show below, it inherits some of the issues of the original surface hopping since the idea of coupled trajectories was not fully “embraced” by all aspects of the algorithm. Instead, our new proposition appears to be more robust and stable, and, thus, we believe that our strategy is an important step forward in the development of novel ideas for nonadiabatic molecular dynamics schemes. It is worth mentioning here that other methodologies exist that are based on coupled trajectories, as for instance full and ab initio multiple spawning (FMS and AIMS), 58 ab initio multiple cloning (AIMC), 59, 60, 61 direct-dynamics variational multi-configurational Gaussian (DD-vMCG) method. 62 In these cases, the coupling among the trajectories emerges from the underlying representation of the nuclear wavepackets associated to the electronic states, that are expressed as linear combinations of moving Gaussians, usually referred to as trajectory basis functions.
To present and to test our new algorithm, the remainder of the paper is structured as follows. First, our ideas of surface-hopping-based algorithms with coupled trajectories derived from the exact factorization of the molecular wavefunction are presented in Section 2. Then, Section 3 assesses the performance of these algorithms on the full-dimensional linear vibronic coupling models for fulvene and for 4-(dimethyloamino)benzonitrile (DMABN), 29 whose molecular counterparts, together with ethylene, have been recently proposed as “molecular Tully models” by Ibele and Curchod. 32 Our conclusions are stated in Section 4.
2 Coupled-trajectory methods combined with surface hopping
The exact factorization (EF) 54, 55 proposes to factor the time-dependent molecular wavefunction $\Psi(\mathbf{r},\mathbf{R},t)$ as the product of a marginal nuclear amplitude $\chi(\mathbf{R},t)$, or wavefunction, and a conditional electronic amplitude $\Phi(\mathbf{r},t;\mathbf{R})$,
$\Psi(\mathbf{r},\mathbf{R},t)=\chi(\mathbf{R},t)\Phi(\mathbf{r},t;\mathbf{R})$ (1)
thus yielding the molecular time-dependent Schrödinger equation (TDSE)
$i\hbar\frac{\partial}{\partial t}\Psi(\mathbf{r},\mathbf{R},t)=\left[\sum_{\nu=1}^{N_{n}}\frac{-\hbar^{2}\boldsymbol{\nabla}_{\nu}^{2}}{2M_{\nu}}+\hat{H}_{el}(\mathbf{r},\mathbf{R})\right]\Psi(\mathbf{r},\mathbf{R},t)$ (2)
in terms of two coupled evolution equations: an effective nuclear TDSE and an electronic evolution equation,
$i\hbar\frac{\partial}{\partial t}\chi(\mathbf{R},t)$ $=\left[\sum_{\nu=1}^{N_{n}}\frac{[-i\hbar\boldsymbol{\nabla}_{\nu}+\mathbf{A}_{\nu}(\mathbf{R},t)]^{2}}{2M_{\nu}}+\epsilon(\mathbf{R},t)\right]\chi(\mathbf{R},t)$ (3)
$i\hbar\frac{\partial}{\partial t}\Phi(\mathbf{r},t;\mathbf{R})$ $=\left[\hat{H}_{el}(\mathbf{r},\mathbf{R})+\hat{U}_{en}[\Phi,\chi](\mathbf{R},t)-\epsilon(\mathbf{R},t)\right]\Phi(\mathbf{r},t;\mathbf{R})$ (4)
In the previous equations, we used the following notation: $\mathbf{r},\mathbf{R}$ indicate the electronic and nuclear configuration space variables, respectively; the molecular Hamiltonian in Eq. (2) is the sum of the nuclear kinetic energy and the electronic Hamiltonian $\hat{H}_{el}(\mathbf{r},\mathbf{R})$; the index $\nu$ labels the nuclei; $M_{\nu}$ is the mass of the nucleus $\nu$.
The nuclear dynamics expressed in Eq. (3) is generated by a time-dependent vector potential $\mathbf{A}_{\nu}(\mathbf{R},t)=\langle\Phi(t;\mathbf{R})|-i\hbar\boldsymbol{\nabla}_{\nu}\Phi(t;\mathbf{R})\rangle_{\mathbf{r}}$ and a time-dependent potential energy surface $\epsilon(\mathbf{R},t)=\langle\Phi(t;\mathbf{R})|\hat{H}_{el}+\hat{U}_{en}-i\hbar\partial_{t}|\Phi(t;\mathbf{R})\rangle_{\mathbf{r}}$, reminiscent of classical electromagnetic potentials and representing the effect of the electrons. Formally, an elegant and rigorous procedure can be introduced to derived classical-like Hamilton’s equations from this effective TDSE (3) to generate nuclear trajectories under the effect of the force determined from $\mathbf{A}_{\nu}(\mathbf{R},t)$ and $\epsilon(\mathbf{R},t)$.63 In turn, the nuclei affect the electronic dynamics entering the evolution equation with their momentum/velocity field and with their spatial distribution encoded in the electron-nuclear coupling operator $\hat{U}_{en}=\sum_{\nu}[-i\boldsymbol{\nabla}_{\nu}-\mathbf{A}_{\nu}(\mathbf{R},t)]^{2}/(2M_{\nu})+(-i\hbar\boldsymbol{\nabla}_{\nu}\chi(\mathbf{R},t)/\chi(\mathbf{R},t)+\mathbf{A}_{\nu}(\mathbf{R},t))\cdot(-i\hbar\boldsymbol{\nabla}_{\nu}-\mathbf{A}_{\nu}(\mathbf{R},t))/M_{\nu}$. Specifically, using the polar representation of the nuclear wavefunction, i.e., expressing it in terms of its modulus $|\chi|$ and phase $S$, we can rewrite the leading 64 term [^1] Leading order in terms of the electron-nuclear mass ratio. of the electron-nuclear coupling operator as
$\frac{-i\hbar\boldsymbol{\nabla}_{\nu}\chi(\mathbf{R},t)}{\chi(\mathbf{R},t)}+\mathbf{A}_{\nu}(\mathbf{R},t)$ $=\left[\boldsymbol{\nabla}_{\nu}S(\mathbf{R},t)+\mathbf{A}_{\nu}(\mathbf{R},t)\right]+i\frac{-\hbar\boldsymbol{\nabla}_{\nu}|\chi(\mathbf{R},t)|^{2}}{2|\chi(\mathbf{R},t)|^{2}}$ (5)
$=\mathbf{P}_{\nu}(\mathbf{R},t)+i\mathbf{Q}_{\nu}(\mathbf{R},t)$ (6)
The term in square brackets on the right-hand side of Eq. (5) defines the nuclear momentum field $\mathbf{P}_{\nu}(\mathbf{R},t)$, including the effect of the time-dependent vector potential, while the quantum momentum $\mathbf{Q}_{\nu}(\mathbf{R},t)$ appears as an imaginary correction to the real momentum field and depends on the spatial nuclear distribution (the second term on the right-hand side of Eq. (5)). If the quantum momentum is neglected in Eq. (6) and in the expression of the electron-nuclear coupling operator, the classical limit of the EF equations leads to Ehrenfest-like equations, 65 and, specifically, the effect of the nuclei on the electronic evolution is simply encoded in the momentum. Equation (4) would then yield a “standard” electronic dynamics as for instance in the original surface hopping 3 scheme. Instead, the presence of the quantum momentum enables to track the delocalization of the nuclear distribution in time and to transfer this information to the electronic dynamics, thus naturally providing a channel for decoherence effects, as discussed abundantly in the literature. 66, 56, 57 Interestingly, the quantum momentum is reminiscent of a coherence-damping term that appears in the evolution equation for the phase-space-dependent elements of the density matrix in QTSH-SDM 52 where auxiliary Gaussians in phase space are introduced.
@@ 2 Coupled-trajectory methods combined with surface hopping
As simple as this, the coupled-trajectory idea emerges in EF: while the momentum field can be interpreted classically, as it is a local property that can be encoded in a trajectory, the very nature of the spatial nuclear distribution is non-local. Therefore, to determine the quantum momentum, the trajectories become coupled since only the whole ensemble yields the necessary non-local information. Extensive analyses of the time-dependent potentials of the EF have justified some additional – necessary for computational efficiency – approximations that ultimately led to the trajectory-based quantum-classical algorithm dubbed coupled-trajectory mixed quantum-classical (CT-MQC). These approximations have been extensively discussed in the literature 67, 68, 19, 69. CT-MQC equations express the electronic evolution and the classical-like nuclear forces in the adiabatic basis, formed by the set of the eigenstates of the electronic Hamiltonian $\hat{H}_{el}$, and it requires to calculate at every time step of dynamics the $3N_{n}$-dimensional nonadiabatic coupling vectors (NACVs) for each pair of electronic states considered in the simulation. In order to circumvent this costly operation, additional simplifications have been considered, namely (i) to use purely adiabatic forces to drive the nuclear dynamics, since the time-dependent scalar potential reproduces piecewise the adiabatic shapes 41, and (ii) to employ the idea of hopping with associated velocity rescaling, to compensate for the effect of the time-dependent vector potential that couples to the nuclear kinetic energy in the nuclear TDSE 42, 70. Based on these observations, the CT-TSH algorithm was proposed 19, 37.
@@
CT-TSH is capable to circumvent the issue of overcoherence via the presence of the quantum momentum in the electronic evolution, which encodes, as argued above, information about the nuclear spatial delocalization. Nonetheless, as we will show below and as previously observed 71, CT-TSH suffers from frustrated hops since, apart from the evaluation of the quantum momentum, the trajectories behave as independent entities and their coupling is not fully exploited, in particular to enforce energy conservation after the hops. Instead, by simply understanding that the trajectories need to cooperate to yield a meaningful approximation of the nuclear quantum dynamics, the additional issues discussed above can be easily cured: if the trajectories share their kinetic energy when hops occur, so as to conserve the total energy of the swarm, frustrated hops are avoided (or strongly reduced to a statistically irrelevant amount) and different velocity rescaling schemes do not affect the outcome of the simulation. Note that the idea of energy conservation over the full ensemble of coupled trajectories has been previously suggested by Villaseco Arribas and Maitra 72, 73 in the context of CT-MQC.
The CT-TSH algorithm 19, 37 is thus summarized with the following equations
$\dot{C}_{I}^{\alpha}(t)=$ $\dot{C}_{I,\textrm{TSH}}^{\alpha}(t)+\dot{C}_{I,\textrm{CT}}^{\alpha}(t)$ (7)
$\mathbf{F}_{\nu}^{\alpha}(t)=$ $-\boldsymbol{\nabla}_{\nu}E_{\textrm{active}}^{\alpha}$ (8)
We used the symbol $C_{I}^{\alpha}(t)$ for the coefficient of the electronic state $I$ in the expansion of the electronic time-dependent conditional amplitude of Eq. (1) in the adiabatic basis. The trajectories are labeled with the index $\alpha$, thus the dependence on $\mathbf{R}\rightarrow\mathbf{R}^{\alpha}(t)$ is simply indicated via the superscript $\alpha$. The electronic coefficients evolve according to a standard “Tully surface hopping” (TSH) term 3
$\dot{C}_{I,\textrm{TSH}}^{\alpha}(t)=-\frac{i}{\hbar}E_{I}^{\alpha}C_{I}^{\alpha}(t)-\sum_{J}\sum_{\nu=1}^{N_{n}}\boldsymbol{v}_{\nu}^{\alpha}(t)\cdot\mathbf{d}_{\nu,IJ}^{\alpha}C_{J}^{\alpha}(t)$ (9)
and to a “coupled-trajectory” (CT) contribution that depends on the quantum momentum
$\dot{C}_{I,\textrm{CT}}^{\alpha}(t)=\sum_{\nu=1}^{N_{n}}\frac{\mathbf{Q}_{\nu}^{\alpha}(t)}{\hbar M_{\nu}}\cdot\left(\mathbf{f}_{\nu,I}^{\alpha}-\sum_{J}\left|C_{J}^{\alpha}(t)\right|^{2}\mathbf{f}_{\nu,J}^{\alpha}\right)C_{I}^{\alpha}(t)$ (10)
In the above equations, we introduced the following symbols: the adiabatic energy of state $I$ at the position $\mathbf{R}^{\alpha}(t)$ as $E_{I}^{\alpha}$, the NACV between states $I$ and $J$ at the position $\mathbf{R}^{\alpha}(t)$ as $\mathbf{d}_{\nu,IJ}^{\alpha}$, the adiabatic force of state $J$ accumulated over time along the trajectory $\mathbf{R}^{\alpha}(t)$ as $\mathbf{f}_{\nu,J}^{\alpha}=\int^{t}(-\boldsymbol{\nabla}_{\nu}E_{J}^{\alpha})d\tau$, the velocity of the nucleus $\nu$ along the trajectory $\mathbf{R}^{\alpha}(t)$ as $\boldsymbol{v}_{\nu}^{\alpha}(t)$. In addition, the classical force in Eq. (8) driving the evolution of the nuclei is purely adiabatic and is given by the gradient of the active state. In CT-TSH, the active state is identified at each time step according to the fewest-switches procedure 3.
In CT-TSH, the coupling among the trajectories is only employed to determine the quantum momentum in $\dot{C}_{I,\textrm{CT}}^{\alpha}(t)$. This is done, as in standard CT-MQC, by reconstructing the nuclear density at each time step as a sum of frozen Gaussians centered at the positions of the trajectories, as abundantly documented in Refs. [74, 75, 19, 68]. The use of frozen Gaussians to reconstruct the nuclear density yields an analytical expression for the quantum momentum. However, this form violates the physical conditions that no population transfer between two electronic states should occur when averaged over the ensemble of trajectories if the NACV between them is zero. This condition needs, thus, to be imposed a posteriori, by introducing a modified expression of the quantum momentum at each time step. 76, 75 Understanding the importance of fully exploiting the coupling among the trajectories in surface hopping leads to the derivation of the combined coupled-trajectory Tully surface hopping (CCT-TSH), that we illustrate below and that we test in Section 3.
CCT-TSH is based on the idea of “energy sharing” to conserve the energy of the swarm of trajectories and to alleviate the frustrated hops. CCT-TSH is based on Eqs. (7) and (8), and we introduce an energy-sharing procedure only at the moment of the hops. We propose three schemes for the energy sharing, which are schematically summarized in Fig. 1: (i) the equity-based scheme, (ii) the overlap-based scheme, and (ii) the quantum-momentum-based scheme. We stress here that, while the idea of energy sharing appears well-founded from a quantum-mechanical viewpoint, its practical implementation in a trajectory-based scheme is quite arbitrary. Therefore, the schemes proposed here are initial attempts to introduce this idea in CCT-TSH and we expect them to evolve based on future additional tests and benchmarks.
@@
Before detailing the energy-sharing schemes, let us introduce the following quantities. The trajectories are labeled with Greek indices, and the hopping trajectory is the trajectory $\alpha$, with kinetic energy indicated with the symbol $E_{\text{kin}}^{\alpha}$. The variation in potential energy for a hop from state $I$ towards a state $J$ of higher energy is $\Delta E^{\alpha}_{\text{hop}}<0$. The Cartesian components of the vectors are indicated with Latin indices, thus for instance the symbol $v_{i}^{\beta}(t)$ stands for the component $i$ of the velocity of the trajectory $\beta$ at time $t$. For scheme (ii), we define the overlap between two trajectories, i.e., $S_{\alpha\beta}$, as the overlap between two Gaussians centered at the positions of the trajectories $\alpha$ and $\beta$. Note that the algorithm does not depend on the value of the width of the Gaussians, as long as they are the same for all trajectories and are large enough to guarantee accounting for all potentially contributing trajectories, since $S_{\alpha\beta}$ is only used to quantify the proximity of the trajectories. In addition, since the trajectories are used to reconstruct the quantum nuclear density in configuration space, we chose to consider their overlap in configuration space and not in phase space (indeed, alternative choices might be equally or even better performing). For scheme (iii), we identify two channels for energy sharing, one driven by the nonadiabatic couplings between states $I$ and $J$, $\boldsymbol{d}_{\nu,IJ}^{\alpha}$, and one driven by the quantum momentum, $\mathbf{Q}_{\nu}^{\alpha}$, both evaluated at the position of the hopping trajectory. Therefore, we use the scalar quantities $\boldsymbol{v}^{\alpha}\cdot\boldsymbol{d}_{IJ}^{\alpha}=\sum_{\nu}\boldsymbol{v}_{\nu}^{\alpha}\cdot\boldsymbol{d}_{\nu,IJ}^{\alpha}$ and $\boldsymbol{q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}=\hbar^{-1}\sum_{\nu}(\mathbf{Q}_{\nu}^{\alpha}/M_{\nu})\cdot(\mathbf{f}_{\nu,I}^{\alpha}-\mathbf{f}_{\nu,J}^{\alpha}))\textrm{Re}[\bar{C}_{J}^{\alpha}(t)C_{I}^{\alpha}(t)]$, to decompose the energy variation to be compensated for the hop to occur. As described above, these scalar quantities appear in the electronic evolution equation of CT-TSH and CCT-TSH, with $\boldsymbol{f}_{\nu,IJ}^{\alpha}=\hbar^{-1}\textrm{Re}[\bar{C}_{J}^{\alpha}(t)C_{I}^{\alpha}(t)](\mathbf{f}_{\nu,I}^{\alpha}-\mathbf{f}_{\nu,J}^{\alpha})$ the coefficient-weighted, time-integrated difference of the adiabatic forces for states $I$ and $J$ and with $\boldsymbol{q}^{\alpha}_{\nu}=\boldsymbol{Q}_{\nu}^{\alpha}/M_{\nu}$ a “pseudo-velocity”. Note that if Eq. (10) is rewritten as
$\dot{C}_{I,\textrm{CT}}^{\alpha}(t)=\sum_{J}\left[\sum_{\nu=1}^{N_{n}}\frac{\mathbf{Q}_{\nu}^{\alpha}(t)}{\hbar M_{\nu}}\cdot\left(\mathbf{f}_{\nu,I}^{\alpha}-\mathbf{f}_{\nu,J}^{\alpha}\right)\bar{C}_{J}^{\alpha}(t)C_{I}^{\alpha}(t)\right]C_{J}^{\alpha}(t)$ (11)
this expression of $\dot{C}_{I,\textrm{CT}}^{\alpha}(t)$ is formally similar to the second term on the right-hand side of $\dot{C}_{I,\textrm{TSH}}^{\alpha}(t)$ in Eq. (9) and justifies the use of the (real part of the) term in square bracket to evaluate the energy sharing.
Using the definitions given above, the energy-sharing schemes are as follows:
@@
    where $v_{i}^{\beta}$$^{\prime}$ is the velocity after rescaling. Weighting the contribution of the trajectories $\beta$ according to their kinetic energy ensures a minimal impact on the trajectory itself. Since the trajectory $\alpha$ uses its entire kinetic energy for the hop, its new velocity after the hop is $\boldsymbol{v}^{\alpha}=0$.
@@
    Quantum-momentum-based. The amount of energy to be rescaled $\Delta E^{\alpha}_{\text{hop}}$ is divided in two contributions: One contribution, $\Delta E^{\alpha}_{\text{NAC}}$, that will be compensated by the kinetic energy of the hopping trajectory $\alpha$ and that will be rescaled along the NACV, and a second contribution, $\Delta E^{\alpha}_{\text{Qmom}}$, that will be compensated by the kinetic energy of the whole swarm of trajectories based on (ii). The two contributions are calculated according to
    $\Delta E^{\alpha}_{\text{NAC}}=\Delta E^{\alpha}_{\text{hop}}\frac{|\boldsymbol{v}^{\alpha}\cdot\boldsymbol{d}_{IJ}^{\alpha}|}{|\boldsymbol{v}^{\alpha}\cdot\boldsymbol{d}_{IJ}^{\alpha}|+|\boldsymbol{q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}|}$ (14)